OpenStudy (foolaroundmath):

Simplify this please: \(\sqrt{\sqrt{5}+2}+\sqrt{\sqrt{5}-2} \)

OpenStudy (shubhamsrg):

sqrt(sqrt(5) +2) = sqrt(5-4) / sqrt(sqrt(5) -2) = 1/ sqrt(sqrt(5) -2) added to sqrt(sqrt(5) -2) we have (1 + sqrt(5) -2) / sqrt(sqrt(5) -2) = (sqrt(5) -1) / sqrt(sqrt(5) -2) also sqrt(sqrt(5) -2) = sqrt(5-4) / sqrt(sqrt(5) +2) substituting, sqrt(sqrt(5) +2) (sqrt(5) -1) lol..maybe not simplified much..hmm

OpenStudy (shubhamsrg):

ohh wait,,leme try another way,, let sqrt(sqrt(5) -2) = b and sqrt(sqrt(5) +2) = a we need to find a+b we know (a+b)^2 = a^2 + b^2 + 2ab substituting, (a+b)^2 = sqrt(5) +2 + sqrt(5) -2 + 2 (sqrt ( 5-4)) =>a+b = sqrt( 2sqrt(5) +2) maybe this helps

OpenStudy (foolaroundmath):

ah nice work .. thank you ..